If personal income taxes are increased by $10, MPC=0.8, then disposable income will _____and consumption will_____.

increase by $10; increase by $8.

stay the same, stay the same

decrease by $10; decrease by $8.

decrease by $10; decrease by $10
Answer #1

Answer

Option 3

decrease by $10; decrease by $8

============

Disposable income =Y-T

Y=income

T=taxes

Disposable income =Y-10

the disposable income decreases by $10

===========

consumption funtion is the form of

C=a+MPC*YD

a=autonomus consumption

YD=disposable income

C=a+0.8*(Y-10)

C=a+0.8Y-8

the consumption decreases by $8